The difference between two numbers is 3. one number Is t. What are the possible values for the other number? Show your Workings.
Answer by Alan3354(69443)

The difference between two numbers is 3. one number Is t. What are the possible values for the other number?
-----------
n = the other number
n = t-3 or
n = 3-t
-------
|n-t| = 3
n = |t-3|
| | is absolute value